OpenStudy (bruno102):

Write the equation of the quadratic function with roots 6 and 10 and a vertex at (8, 2).

OpenStudy (anonymous):

start with \[y=a(x-6)(x-10)\] then find \(a\) be replacing \(x\) by \(8\) and \(y\) by \(2\)

OpenStudy (bruno102):

so 2=a(8-6)(8-10)

OpenStudy (anonymous):

yes

OpenStudy (bruno102):

and isn't that 2=a(2)(-2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what is \(a\) ?

OpenStudy (bruno102):

well 2=a(-4) so a= -1/2?

OpenStudy (anonymous):

yes

OpenStudy (bruno102):

Oh ok, so how would I rewrite the equation?

OpenStudy (anonymous):

\[y=-\frac{1}{2}(x-6)(x-10)\] and then multiply out i would multiply \((x-6)(x-10)\) first, then divide each coefficient by \(-2\)

OpenStudy (bruno102):

Thank you!

OpenStudy (anonymous):

yw